OpenStudy (laddiusmaximus):

estimate instantaneous rate of change at the point x=5 for f(x)=lnx

OpenStudy (amistre64):

the derivative of lnx at x=5

OpenStudy (amistre64):

1/5 maybe?

OpenStudy (laddiusmaximus):

That is not what it is asking. Im supposed to use the average velocity formula to figure it out and its not working

OpenStudy (amistre64):

the average velocity is for an average rate of change; your spose to be using the limit of that formula as h goes to 0 or some other denominator

OpenStudy (laddiusmaximus):

I know that but we havent gotten that far. We are only using rates of change.

OpenStudy (amistre64):

instantaneous velocity is the derivatie; which can be defined as:\[\lim_{h->0}\frac{ln(x+h)-ln(x)}{h }\]

OpenStudy (amistre64):

then use the avg stuff with small changes

OpenStudy (amistre64):

ln(5)-ln(5.0001)/.0001

OpenStudy (amistre64):

or something akin to it

OpenStudy (laddiusmaximus):

the way the book and cramster teaches it is way different. im supposed to use numbers a little more and a little less than five

OpenStudy (laddiusmaximus):

like 4.9 and 5.01

OpenStudy (amistre64):

teh smaller the change, the more exact the estimate

OpenStudy (laddiusmaximus):

(5)^3 +5-((4.9)^3 +4.9)/ 5-4.9

OpenStudy (laddiusmaximus):

well it asks for estimates and all the ones I give are wrong

OpenStudy (amistre64):

cubed stuff?

OpenStudy (amistre64):

you said lnx

OpenStudy (laddiusmaximus):

I end up with 68.75, 73.04,74.51,75.85

OpenStudy (laddiusmaximus):

I dont know man. Ive been beating my head against the wall for the last two hours and the book is nigh on useless

OpenStudy (amistre64):

ln(5.00001/5)/.00001 = .2000...... approx 1/5 http://www.wolframalpha.com/input/?i=ln%285.00001%2F5%29%2F.00001

OpenStudy (laddiusmaximus):

it says that the instantaneous rate of change is estimated by computing the average rate of smaller intervals

OpenStudy (amistre64):

thats what i did

OpenStudy (laddiusmaximus):

well what am I doing then? because im lost. I dont understand how to set up the problem.

OpenStudy (amistre64):

ln(5.00001) is a small change from ln(5) right?

OpenStudy (amistre64):

in fact is .00001 small of a change

OpenStudy (laddiusmaximus):

ok i follow that. but what then? how do I set up the problem to solve it? so its [5, 5.001]?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

or 4.99999 to 5 but its easier to see the change when it grows a little bigger

OpenStudy (laddiusmaximus):

how do i plug it in? i see a formula for average rate of change, but it doesnt give one for instantaneous

OpenStudy (amistre64):

im not sure what it is youre looking at ....

OpenStudy (amistre64):

if its about inputing into a program, i got no idea what it wants then

OpenStudy (amistre64):

the instantaneous rate of change at ln(5) = 1/5 = .2

OpenStudy (amistre64):

so how you got those other vlaues I cant say

OpenStudy (laddiusmaximus):

im sorry. the book does a very crappy job of explaining things